# Stick-breaking Attention

### v1

@@ Abstract
The self-attention mechanism traditionally relies on the softmax operator, necessitating positional embeddings like RoPE, or position biases to account for token order. But current methods using still face length generalisation challenges. We propose an alternative attention mechanism based on the stick-breaking process: For each token before the current, we determine a break point $\beta_{i,j}$, which represents the proportion of the remaining stick to allocate to the current token. We repeat the process until the stick is fully allocated, resulting in a sequence of attention weights. This process naturally incorporates recency bias, which has linguistic motivations for grammar parsing (Shen et al., 2017). We study the implications of replacing the conventional softmax-based attention mechanism with stick-breaking attention. We then discuss implementation of numerically stable stick-breaking attention and adapt Flash Attention to accommodate this mechanism. When used as a drop-in replacement for current softmax+RoPE attention systems, we find that stick-breaking attention performs competitively with current methods on length generalisation and downstream tasks. Stick-breaking also performs well at length generalisation, allowing a model trained with $2^{11}$ context window to perform well at $2^{14}$ with perplexity improvements.
@@ 1 Introduction
The Transformer architecture (Vaswani et al., 2017) uses a self-attention mechanism based on the softmax operator that enables the model to weigh the importance of different tokens in the input data. However, the reliance on softmax requires using positional embeddings to introduce information about the order of tokens, as the attention mechanism itself is permutation-invariant. The sinusoidal position embedding as proposed in Vaswani et al. (2017) has since evolved to relative positional embeddings (Shaw et al., 2018). Learned relative positional biases were used in the T5 model (Raffel et al., 2020), and later fixed relative positional biases in Press et al. (2021). At the time of this writing, a commonly used form of position embedding is RoPE (Su et al., 2021). Allen-Zhu & Li (2023) observe that, in a context-free grammar parsing setting, attention mechanisms attend to the “most adjacent” non-terminal. This suggests an inclination to attend to the most recent entry that matches a given criteria. However, even with relative position information, it is possible to overfit on specific relative positions, resulting in failure to generalise. Kazemnejad et al. (2024) show that decoder-only Transformers with No Positional Embeddings (NoPE) can implicitly recover positional information, experimental results suggest that NoPE Transformers generalise better on length. While this is promising, a higher attention score from an irrelevant token in the sequence can function as a distractor (Kazemnejad et al., 2024; Xiao et al., 2023).
The stick-breaking process may have properties that can alleviate the previously mentioned issues, and possess the ‘most recent’ bias from Allen-Zhu & Li (2023) we want. For a token at position $j$ attending to position $i$, suppose the attention weight is given by:
@@
Logits $z_{i,j}=\frac{\displaystyle{\bm{q}}_{i}^{\top}{\bm{k}}_{j}}{\displaystyle\sqrt{{d_{\mathrm{head}}}}}$
Softmax ${\bm{A}}_{i,j}=\frac{\displaystyle\exp(z_{i,j})}{\displaystyle\sum_{j^{\prime}\leq i}\exp(z_{i,j^{\prime}})}$
@@
Figure 1: Differences in formulation between stick-breaking and softmax. Stick-breaking assigns high weights to the most recent high logit, while softmax will assign equal weightage to equal logits.
Shen et al. (2017) makes a similar observation as in Allen-Zhu & Li (2023), and explicitly uses a stick-breaking process to model local structure. Csordás et al. (2021) introduces Geometric attention, named after the Geometric distribution, which only has one parameter $p$, which gives the probability of success per trial. The geometric distribution then gives the probability for which $k$ trials are needed for the first success: $(1-p)^{k-1}p$. But in stick-breaking and in Geometric attention (Csordás et al., 2021), each $p$ is assigned a different value that corresponds to the attention score between two tokens.
@@ 2 Stick-breaking Attention
${\bm{o}}_{j}$ $=\sum_{i=1}^{j-1}{\bm{A}}_{i,j}\cdot{\bm{v}}_{i},$ ${\bm{A}}_{i,j}$ $=\beta_{i,j}\prod_{i<k<j}\left(1-\beta_{k,j}\right),$ $\beta_{i,j}$ $=\sigma\left(z_{i,j}\right),$ $z_{i,j}$ $=\frac{{\bm{q}}_{j}^{\top}{\bm{k}}_{i}}{\sqrt{{d_{\mathrm{head}}}}}$ (1)
Equation 1 is the main difference between our proposal and softmax attention. As discussed earlier and in Csordás et al. (2021), this parameterisation biases towards recency. Specifically, for any pair of $i$ and $i^{\prime}$ such that $i-j<i^{\prime}-j$ and $z_{i,j}=z_{i,j}$, then ${\bm{A}}_{i,j}\geq{\bm{A}}_{i^{\prime},j}$. Consequently, this imposes an ordering on the attention, and we do not use position embeddings with the query and key embeddings.
We consider two sets of logits, $z_{i,j}$ and $z^{\prime}_{i,j}$ and their respective attention weights ${\bm{A}}_{i,j}$ and ${\bm{A}}^{\prime}_{i,j}$. If $z_{i,k}=z^{\prime}_{i,k}$ for $i<k<j$, then ${\bm{A}}_{i,j}={\bm{A}}^{\prime}_{i,j}$. Further, if $\sum_{k=i}^{j-1}{\bm{A}}_{k,j}=1$, then the output is invariant to appending additional context earlier than $i$. This means that unlike softmax attention, a high attention score further back in the sequence does not ‘distract’ from a more recent high $z_{i,j}$ score. We note that $\sum_{i=1}^{j-1}{\bm{A}}_{i,j}\leq 1$, which allows this attention mechanism to attend to nothing when all $\beta_{i,j}=0$.
@@ 3 Related Work
Each stick-breaking attention head at every time-step can be viewed as the hidden state of the final step of a selective State-space Model (SSM;Gu & Dao 2023). For a given time-step $j$, consider the following SSM and its convolutional form (as described in Merrill et al. 2024):
$\hat{{\bm{o}}}_{i,j}=(1-\beta_{i,j})\cdot\hat{{\bm{o}}}_{i-1}+\beta_{i,j}\cdot{\bm{v}}_{j},$ ${\bm{o}}_{j}=\hat{{\bm{o}}}_{j-1,j}=\sum_{1\leq i\leq j}\left(\beta_{i}\prod_{i<k<j}(1-\beta_{k})\right)\cdot{\bm{v}}_{i},$ (2)
which is equivalent to the first term in Equation 1. Typically, an attention layer for a Transformer with hidden dimension $d_{\mathrm{hidden}}$ has $h$ heads such that $d_{\mathrm{hidden}}=h\cdot{d_{\mathrm{head}}}$. For equivalence with an SSM, we need a constant query vector and each dimension as a separate head, e.g. ${\bm{q}}_{i}=\mathbf{1}$ for all $i$, and $h=d_{\mathrm{hidden}},{d_{\mathrm{head}}}=1$.
@@ 4 Implementation
Implementing stick-breaking attention naively in PyTorch results in realising the $L^{2}$ matrix for the attention logits (where $L$ is the length of the input). FlashAttention (Dao et al., 2022) reduces the memory footprint of attention by side-stepping the $O(L^{2})$ memory complexity of realising the attention matrix. In order to achieve this, it only realises tiles of the attention logits and weights at a time, and accumulates the resulting weighted sum of ${\bm{v}}_{i}$. In this section we detail the important differences between computing softmax attention and stick-breaking.
(a) Forward pass
(b) 2-phase backward pass
(c) Block Skipping
Figure 2: Thread tile assignments for a given attention head and a sequence. Tiles coloured the same are processed by the same thread. For stick-breaking, the forward pass has to be computed from right-to-left, while the backward pass is computed in two-phases: left-to-right for $\nabla{\bm{Q}}$, and top-to-bottom for $\nabla{\bm{V}},\nabla{\bm{K}}$. Uncoloured tiles are not computed: upper right tiles are not used in causal language modelling, and in the case of block skipping, some blocks can be skipped if all entries have summed to 1. The accumulated values ${\color[rgb]{0.65234375,0.109375,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.65234375,0.109375,0}{\bm{a}}},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\bm{b}}}$ are highlighted at the borders for a visual representation of where they are used.
@@
Where $\log\left(1+\exp(\cdot)\right)$ is commonly known as the $\mathrm{softplus}$ operation, which we further numerically stabilise with the following computation:
@@
to prevent overflowing of $\exp(x)$.
@@
The above equations dictate the direction of the order of computation for our implementation. For the forward pass (Eqn. 3), we compute from $j$ to $1$, backwards through time and accumulate $\sum_{k=i+1}^{j}\log\left(1+\exp(z_{k,j})\right)$. For backward pass (Eqn. 5), we compute from $1$ to $j$, accumulating $\sum_{j^{\prime}=1}^{j-1}\frac{\partial\mathcal{L}}{\partial\tilde{{\bm{A}}_{i,j^{\prime}}}}$.
@@ 4.1 Triton Specifics
We modify the Triton implementation of Flash Attention for accelerating the stick-breaking attention. In theory, the memory complexity of implementing stick-breaking attention would be similar to that of Flash Attention, $O(L)$. Due to the limitations of Triton, we borrow from Dao (2023) and use a two-phased backward pass: the first phase accumulating gradients for ${\bm{Q}}$ , and the second phase accumulates gradients for ${\bm{K}},{\bm{V}}$. Due to the intermediate accumulated values ${\color[rgb]{0.65234375,0.109375,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.65234375,0.109375,0}{\bm{a}}},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\bm{b}}}$, our implementation results in a $O(L^{2})$ memory complexity. However, the constant factor in the quadratic complexity here is inverse of the predetermined block size $d_{\mathrm{block}}$: larger block sizes reduces the memory allocation. In our implementation, $d_{\mathrm{block}}=64$. Algorithms 1 details the forward pass (illustrated in Figure 2(a)), and Algorithms 2 & 3 details the backward pass (illustrated in Figure 2(b)).
@@
  ${\bm{Q}}\leftarrow$ load block $i$ of ${\bm{Q}}$ // load High Bandwidth Memory (HBM) to Static RAM (SRAM), ${\bm{Q}}:d_{\mathrm{block}}\times{d_{\mathrm{head}}}$
@@
     ${\bm{V}}\leftarrow$ load block $k$ of ${\bm{V}}$ // ${\bm{V}}:d_{\mathrm{block}}\times{d_{\mathrm{head}}}$
@@
[…]
@@
     ${\color[rgb]{0.65234375,0.109375,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.65234375,0.109375,0}{\bm{a}}}\rightarrow$ store block $(i,k)$ of ${\color[rgb]{0.65234375,0.109375,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.65234375,0.109375,0}{\bm{M}}}$ // store from SRAM to HBM
@@
  ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\bm{b}}}\leftarrow\mathbf{0},{\bm{O}}\leftarrow\mathbf{0}$
  $\nabla{\bm{O}}\leftarrow$ load block $i$ of $\nabla{\bm{O}}$
  ${\bm{Q}}\leftarrow$ load block $i$ of ${\bm{Q}}$
@@
     ${\bm{K}}\leftarrow$ load block $k$ of ${\bm{K}}$
     ${\bm{V}}\leftarrow$ load block $k$ of ${\bm{V}}$
     ${\color[rgb]{0.65234375,0.109375,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.65234375,0.109375,0}{\bm{a}}}\leftarrow$ load block $(i,k)$ of ${\color[rgb]{0.65234375,0.109375,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.65234375,0.109375,0}{\bm{M}}}$
@@
     $\nabla\tilde{{\bm{A}}}\leftarrow{\bm{A}}\odot(\nabla{\bm{O}}{\bm{V}}^{\top})$
     ${\nabla\tilde{{\bm{A}}}}_{\mathrm{cs}}\leftarrow\mathrm{cumsum}_{\rightarrow}\nabla\tilde{{\bm{A}}}$
     $\nabla{\bm{Z}}\leftarrow\nabla\tilde{{\bm{A}}}-(1-\exp({\bm{L}}))\odot\left({\nabla\tilde{{\bm{A}}}}_{\mathrm{cs}}+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\bm{b}}}\right)$
     ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\bm{b}}}\leftarrow{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\bm{b}}}+\sum_{\rightarrow}\nabla\tilde{{\bm{A}}}$
     ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\bm{b}}}\rightarrow$ store block $(i,k)$ of ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\bm{N}}}$
@@
Algorithm 2 Backward-${\bm{q}}$ thread $i$
  $\nabla{\bm{K}}\leftarrow\mathbf{0},\nabla{\bm{V}}\leftarrow\mathbf{0}$
  ${\bm{K}}\leftarrow$ load block $j$ of ${\bm{k}}$
  ${\bm{V}}\leftarrow$ load block $j$ of ${\bm{v}}$
  for $k$ in $j\ldots J$ do
     ${\bm{Q}}\leftarrow$ load block $k$ of ${\bm{q}}$
     $\nabla{\bm{O}}\leftarrow$ load block $k$ of $\nabla{\bm{o}}$
     ${\color[rgb]{0.65234375,0.109375,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.65234375,0.109375,0}{\bm{a}}}\leftarrow$ load block $(k,j)$ of ${\color[rgb]{0.65234375,0.109375,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.65234375,0.109375,0}{\bm{M}}}$
     ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\bm{b}}}\leftarrow$ load block $(k,j)$ of ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\bm{N}}}$
     Do 1
     $\nabla\tilde{{\bm{A}}}\leftarrow{\bm{A}}\odot(\nabla{\bm{O}}{\bm{V}}^{\top})$
     ${\nabla\tilde{{\bm{A}}}}_{\mathrm{cs}}\leftarrow\mathrm{cumsum}_{\rightarrow}\nabla\tilde{{\bm{A}}}$
     $\nabla{\bm{Z}}\leftarrow\nabla\tilde{{\bm{A}}}-(1-\exp({\bm{L}}))\odot\left({\nabla\tilde{{\bm{A}}}}_{\mathrm{cs}}+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\bm{b}}}\right)$
     $\nabla{\bm{K}}\leftarrow\nabla{\bm{K}}+{\bm{A}}^{\top}\nabla{\bm{O}}$
     $\nabla{\bm{V}}\leftarrow\nabla{\bm{V}}+\nabla{\bm{Z}}^{\top}{\bm{Q}}$
  end for
  $\nabla{\bm{K}}\rightarrow$ store block $j$ of $\nabla{\bm{K}}$
  $\nabla{\bm{V}}\rightarrow$ store block $j$ of $\nabla{\bm{V}}$
Algorithm 3 Backward-${\bm{v}}$,${\bm{k}}$ thread $j$
@@
We measure throughput on Dolomite Engine (Mishra, 2024) with a 1B class model on a node with 8 H100 GPUs. Stick-breaking attains a throughput of 16.27 billion tokens / day. This is a a 19.9% performance drop compared to Flash Attention, which attains throughput of 19.5 billion tokens / day. Despite the slowdown, the Triton implementation allows the method to be used for long sequences as the naive Torch implementation causes out-of-memory issues.

### v2

# Scaling Stick-Breaking Attention: An Efficient Implementation and In-depth Study

@@ Abstract
The self-attention mechanism traditionally relies on the softmax operator, necessitating positional embeddings like RoPE, or position biases to account for token order. But current methods using still face length generalisation challenges. We investigate an alternative attention mechanism based on the stick-breaking process in larger scale settings. The method works as follows: For each token before the current, we determine a break point, which represents the proportion of the stick, the weight of the attention, to allocate to the current token. We repeat this on the remaining stick, until all tokens are allocated a weight, resulting in a sequence of attention weights. This process naturally incorporates recency bias, which has linguistic motivations for grammar parsing (Shen et al., 2017). We study the implications of replacing the conventional softmax-based attention mechanism with stick-breaking attention. We then discuss implementation of numerically stable stick-breaking attention and adapt Flash Attention to accommodate this mechanism. When used as a drop-in replacement for current softmax+RoPE attention systems, we find that stick-breaking attention performs competitively with current methods on length generalisation and downstream tasks. Stick-breaking also performs well at length generalisation, allowing a model trained with $2^{11}$ context window to perform well at $2^{14}$ with perplexity improvements.
@@ 1 Introduction
The Transformer architecture (Vaswani et al., 2017) uses a self-attention mechanism based on the softmax operator that enables the model to weigh the importance of different tokens in the input data. However, the reliance on softmax requires using positional embeddings to introduce information about the order of tokens, as the attention mechanism itself is permutation-invariant. The sinusoidal position embedding as proposed in Vaswani et al. (2017) has since evolved to relative positional embeddings (Shaw et al., 2018). Learned relative positional biases were used in the T5 model (Raffel et al., 2020), and later fixed relative positional biases in Press et al. (2021). At the time of writing, a commonly used form of position embedding is RoPE (Su et al., 2021). Allen-Zhu & Li (2023) observe that, in a context-free grammar parsing setting, attention mechanisms attend to the “most adjacent” non-terminal. This suggests an inclination to attend to the most recent entry that matches a given criteria. However, even with relative position information, it is possible to overfit on specific relative positions, resulting in failure to generalise. Kazemnejad et al. (2024) show that decoder-only Transformers with No Positional Embeddings (NoPE) can implicitly recover positional information, experimental results suggest that NoPE Transformers generalise better on length. While this is promising, a higher attention score from an irrelevant token in the sequence can function as a distractor (Kazemnejad et al., 2024; Xiao et al., 2023).
The stick-breaking process may have properties that can alleviate the previously mentioned issues, and possess the ‘most recent’ bias from Allen-Zhu & Li (2023) that we want. For a token at position $j$ attending to position $i$, suppose the attention weight is given by:
@@
Logits $z_{i,j}=\frac{\displaystyle{\bm{q}}_{j}^{\top}{\bm{k}}_{i}}{\displaystyle\sqrt{{d_{\mathrm{head}}}}}$
Softmax ${\bm{A}}_{i,j}=\displaystyle\frac{\exp(z_{i,j})}{\sum_{k=1}^{j}\exp(z_{k,j})}$
@@
Figure 1: Differences in formulation between stick-breaking and softmax. Stick-breaking assigns high weights to the most recent high logit, while softmax will assign equal weightage to equal logits. $\sigma(\cdot)$ can be any function $\mathbb{R}\rightarrow(0,1)$. In this paper we use $\sigma(x)=\frac{1}{1+exp(-x)}$
Shen et al. (2017) makes a similar observation as in Allen-Zhu & Li (2023), and explicitly uses a stick-breaking process to model local structure. Csordás et al. (2021) has a similar construction which they call Geometric attention after the Geometric distribution. Geometric distribution only has one parameter $p$, which gives the probability of success per trial. The geometric distribution then gives the probability for which $k$ trials are needed for the first success: $(1-p)^{k-1}p$. But in both stick-breaking and in Geometric attention (Csordás et al., 2021), each $p$ is assigned a different value that corresponds to the attention score between two tokens.
@@ 2 Stick-breaking Attention
${\bm{o}}_{j}$ $=\sum_{i=1}^{j-1}{\bm{A}}_{i,j}{\bm{v}}_{i},$ ${\bm{A}}_{i,j}$ $=\beta_{i,j}\prod_{i<k<j}\left(1-\beta_{k,j}\right),$ $\beta_{i,j}$ $=\sigma\left(z_{i,j}\right),$ $z_{i,j}$ $=\frac{{\bm{q}}_{j}^{\top}{\bm{k}}_{i}}{\sqrt{{d_{\mathrm{head}}}}}$ (1)
Equation 1 is the main difference between our proposal and softmax attention. As discussed earlier and in Csordás et al. (2021), this parameterisation biases towards recency. Specifically, for any pair of $i$ and $i^{\prime}$ such that $|j-i|<|j-i^{\prime}|$ and $z_{i,j}=z_{i^{\prime},j}$, then ${\bm{A}}_{i,j}\geq{\bm{A}}_{i^{\prime},j}$. Consequently, this imposes an ordering on the attention, and we do not use position embeddings with the query and key embeddings.
We consider two sets of logits, $z_{i,j}$ and $z^{\prime}_{i,j}$ and their respective attention weights ${\bm{A}}_{i,j}$ and ${\bm{A}}^{\prime}_{i,j}$. If $z_{i,k}=z^{\prime}_{i,k}$ for $i<k<j$, then ${\bm{A}}_{i,j}={\bm{A}}^{\prime}_{i,j}$. This means that unlike softmax attention, a high attention score further back in the sequence does not ‘distract’ from a more recent high $z_{i,j}$ score. Further, if $\sum_{k=i}^{j-1}{\bm{A}}_{k,j}=1$, then the output is invariant to appending additional context earlier than $i$. We note that $\sum_{i=1}^{j-1}{\bm{A}}_{i,j}\leq 1$, which allows this attention mechanism to attend to nothing when all $\beta_{i,j}=0$. We discuss a strategy to deal with the remaining attention weight in Appendix LABEL:sec:remainder.
@@ 3 Related Work
Each stick-breaking attention head at every time-step can be viewed as the hidden state of the final step of a selective State-space Model (SSM;Gu & Dao 2023). For a given time-step $j$, consider the following SSM in its recurrent form and its corresponding convolutional form (as described in Merrill et al. 2024):
$\textrm{Let}\qquad\hat{{\bm{o}}}_{i,j}$ $=(1-\beta_{i,j})\cdot\hat{{\bm{o}}}_{i-1,j}+\beta_{i,j}\cdot{\bm{v}}_{j},$ (2)
$\textrm{then}\qquad{\bm{o}}_{j}$ $=\hat{{\bm{o}}}_{j-1,j}=\sum_{1\leq i\leq j}\left(\beta_{i,j}\prod_{i<k<j}(1-\beta_{k,j})\right)\cdot{\bm{v}}_{i},$ (3)
which is equivalent to the first term in Equation 1. Intuitively, this means that the output head of stick-breaking attention is equivalent to the end state of a single-gate selective SSM. Typically, an attention layer for a Transformer with hidden dimension $d_{\mathrm{hidden}}$ has $h$ heads such that $d_{\mathrm{hidden}}=h\cdot{d_{\mathrm{head}}}$. For equivalence with an SSM, we need a constant query vector and each dimension as a separate head, e.g. ${\bm{q}}_{i}=\mathbf{1}$ for all $i$, and $h=d_{\mathrm{hidden}},{d_{\mathrm{head}}}=1$.
@@ 4 Implementation
Implementing stick-breaking attention naively in PyTorch results in realising the $L^{2}$ matrix for the attention logits (where $L$ is the length of the input). FlashAttention (Dao et al., 2022) reduces the memory footprint of attention by side-stepping the $O(L^{2})$ memory complexity of realising the attention matrix. In order to achieve this, it only realises blocks of the attention weights during computation, and accumulates the resulting weighted sum of ${\bm{v}}_{i}$. We take a similar approach to speeding up stick-breaking attention, allowing it to be used for longer sequences.
@@
Where $\log\left(1+\exp(\cdot)\right)$ is commonly known as the $\mathrm{softplus}$ operation. See Appendix A for the full derivation. We further numerically stabilise $\mathrm{softplus}$ with the following computation:
@@
to prevent overflowing of $\exp(x)$. To further speed up the computation of $\mathrm{softplus}$, we also write the operation in Parallel Thread Execution (PTX).
@@
The above equations dictate the direction of the order of computation for our implementation. For the forward pass (Eqn. 4), we compute from $j$ to $1$, backwards through time and accumulate $\sum_{k=i+1}^{j}\log\left(1+\exp(z_{k,j})\right)$. For backward pass (Eqn. 6), we compute from $1$ to $j$, accumulating $\sum_{j^{\prime}=1}^{j-1}\frac{\partial\mathcal{L}}{\partial\tilde{{\bm{A}}_{i,j^{\prime}}}}$.
@@ 4.1 Triton Specifics
${\bm{Q}}$${\bm{K}},{\bm{V}}$${\bm{O}}$Accumulate ${\bm{O}}$
(a) Forward pass
$\nabla{\bm{Q}}$$\nabla{\bm{K}},\nabla{\bm{V}}$$\nabla{\bm{O}}$Accumulate $\nabla{\bm{Q}}$atomic add
(b) Backward pass
$\nabla{\bm{Q}}$$\nabla{\bm{K}},\nabla{\bm{V}}$$\nabla{\bm{O}}$
(c) Block skipping
Figure 2: Thread tile assignments for a given attention head and a sequence. Tiles coloured the same are processed by the same thread. For stick-breaking, the forward pass has to be computed from right-to-left, while the backward pass is computed in the reverese order. Uncoloured tiles are not computed: upper right tiles are not used in causal language modelling, and in the case of block skipping, some blocks can be skipped if all entries have summed to 1.
We modify the Triton implementation of Flash Attention for accelerating the stick-breaking attention. We used a similar approach as in Flash Attention 2 (Dao 2023; FA2), loading blocks of ${\bm{Q}}$ from High Bandwidth Memory (HBM) into Static random-access memory (SRAM). In the forward pass, we accumulate towards the start of the sequence, loading blocks of ${\bm{K}}$ and ${\bm{V}}$ the values ${\color[rgb]{0.65234375,0.109375,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.65234375,0.109375,0}{\bm{a}}}$, finally writing ${\color[rgb]{0.65234375,0.109375,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.65234375,0.109375,0}{\bm{a}}}$ and ${\bm{O}}$ as the output into HBM. In the backward pass, we differ from the FA2 approach. FA2 accumulates $\nabla{\bm{K}},\nabla{\bm{V}}$, and makes atomic adds toward $\nabla{\bm{Q}}$, which allows for fewer synchronisations during the computation of the backward pass. However, in the case of stick-breaking, the accumulation of the gradients have to be in the reverse direction of the cumulative sum. As a result, we have to do atomic adds for both $\nabla{\bm{K}}$ and $\nabla{\bm{V}}$. The resulting memory complexity of our approach is still $O(L)$, which allows us to train large models with stick-breaking.
In our implementation, the block size is $d_{\mathrm{block}}=64$. Algorithms 1 details the forward pass (illustrated in Figure 2(a)), and Algorithm 2 details the backward pass (illustrated in Figure 2(b)).
@@
  ${\bm{Q}}\leftarrow$ load block $i$ of ${\bm{Q}}$ // load from HBM to SRAM, ${\bm{Q}}:d_{\mathrm{block}}\times{d_{\mathrm{head}}}$
@@
     ${\bm{V}}\leftarrow$ load block $k$ of ${\bm{V}}$
[…]
@@
     ${\bm{L}}\leftarrow-\mathrm{softplus}({\bm{Z}})$ // ${\bm{L}}:d_{\mathrm{block}}\times d_{\mathrm{block}}$
@@
  ${\color[rgb]{0.65234375,0.109375,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.65234375,0.109375,0}{\bm{a}}}\rightarrow$ store block $i$ of ${\color[rgb]{0.65234375,0.109375,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.65234375,0.109375,0}{\bm{M}}}$ // store from SRAM to HBM
@@
  ${\bm{b}}\leftarrow\mathbf{0}$ // init ${\bm{b}}:d_{\mathrm{block}}\times 1$
  $\nabla{\bm{O}}\leftarrow$ load block $i$ of $\nabla{\bm{O}}$ // $\nabla{\bm{O}}:d_{\mathrm{block}}\times{d_{\mathrm{head}}}$
  ${\bm{Q}}\leftarrow$ load block $i$ of ${\bm{Q}}$ // ${\bm{Q}}:d_{\mathrm{block}}\times{d_{\mathrm{head}}}$
  ${\color[rgb]{0.65234375,0.109375,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.65234375,0.109375,0}{\bm{a}}}\leftarrow$ load block $i$ of ${\color[rgb]{0.65234375,0.109375,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.65234375,0.109375,0}{\bm{M}}}$ // ${\color[rgb]{0.65234375,0.109375,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.65234375,0.109375,0}{\bm{a}}}:d_{\mathrm{block}}\times 1$
@@
     ${\color[rgb]{0.65234375,0.109375,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.65234375,0.109375,0}{\bm{a}}}\leftarrow{\color[rgb]{0.65234375,0.109375,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.65234375,0.109375,0}{\bm{a}}}-\sum_{\leftarrow}{\bm{L}}$
     ${\bm{A}}\leftarrow\exp\left({\bm{Z}}+\mathrm{cumsum}_{\leftarrow}{\bm{L}}+{\color[rgb]{0.65234375,0.109375,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.65234375,0.109375,0}{\bm{a}}}\right)$
     $\nabla\tilde{{\bm{A}}}\leftarrow{\bm{A}}\odot(\nabla{\bm{O}}{\bm{V}}^{\top})$ // $\nabla\tilde{{\bm{A}}}:d_{\mathrm{block}}\times d_{\mathrm{block}}$
     $\nabla{\bm{Z}}\leftarrow\nabla\tilde{{\bm{A}}}-(1-\exp({\bm{L}}))\odot\left(\mathrm{cumsum}_{\rightarrow}\nabla\tilde{{\bm{A}}}+{\bm{b}}\right)$ // $\nabla{\bm{Z}}:d_{\mathrm{block}}\times d_{\mathrm{block}}$
     ${\bm{b}}\leftarrow{\bm{b}}+\sum_{\rightarrow}\nabla\tilde{{\bm{A}}}$
@@
     ${\bm{A}}^{\top}\nabla{\bm{O}}\rightarrow\textrm{atomic add block $k$~{}}\nabla{\bm{K}}$ // accumulate gradient for $\nabla{\bm{K}}$ and $\nabla{\bm{V}}$
     $\nabla{\bm{Z}}^{\top}{\bm{Q}}\rightarrow\textrm{atomic add block $k$~{}}\nabla{\bm{V}}$
@@
Algorithm 2 Backward thread $i$
@@
We measure throughput on Dolomite Engine (Mishra, 2024) with a 1B class model on a node with 8 H100 GPUs. Stick-breaking attains a throughput of 21 billion tokens / day. This is a 29% performance drop compared to FA2, which attains a throughput of 29.5 billion tokens / day. Despite the slowdown, the Triton implementation allows the method to be used for long sequences as the naive Torch implementation causes out-of-memory issues. Further optimisation of the kernel may be possible if written in CUDA, taking advantage of synchronisation features the language exposes in the hardware.